# What Did I Do Wrong? Quantifying LLMs’ Sensitivity and Consistency to Prompt Engineering

### v1

@@ Abstract
Large Language Models (LLMs) changed the way we design and interact with software systems. Their ability to process and extract information from text has drastically improved productivity in a number of routine tasks. Developers that want to include these models in their software stack, however, face a dreadful challenge: debugging their inconsistent behavior across minor variations of the prompt. We therefore introduce two metrics for classification tasks, namely sensitivity and consistency, which are complementary to task performance. First, sensitivity measures changes of predictions across rephrasings of the prompt, and does not require access to ground truth labels. Instead, consistency measures how predictions vary across rephrasings for elements of the same class. We perform an empirical comparison of these metrics on text classification tasks, using them as guideline for understanding failure modes of the LLM. Our hope is that sensitivity and consistency will be powerful allies in automatic prompt engineering frameworks to obtain LLMs that balance robustness with performance.
@@ 1 Introduction
This famous quote refers to the innate tendency of computer scientists to choose poor names for a program’s variables despite the existence of good coding practices (McConnell, 2004). From a purely practical perspective, this is not a problem as long as the program does its job, but can we still argue the same in the era of Large Language Models?
@@
Figure 1: Real example of the behavior of GPT3.5 when classifying a question in terms of what it is referring to. A slight change in the definition of the class “ENTY” causes a minor prompt variation that disrupts the LLM’s prediction. This happens under the hood, making it very hard for a developer to debug the program. Note that the same might happen, for instance, if the ordering of the labels or the fields’ names is changed.
From a practical standpoint, integrating LLMs in a software infrastructure introduces additional complexities, from choosing good prompt engineering strategies to parsing and controlling the output format of responses. In an effort to simplify the LLM into a straightforward function call, new software libraries like Instructor (Liu, 2024) have emerged, reducing even further the entry barrier of LLMs for the ordinary programmer. Instructor requires to define target labels as static fields of a Python class, which is then automatically converted into a standardized JSON schema. The schema and the input sentence are then used to produce an LLM response via a mechanism known as Function Calling (Kang et al., 2023). Figure 1 provides a real example of text classification using Instructor. Here, the assumption of the programmer should be that small changes to the labels’ descriptions, as well as the labels’ ordering in the code, should not affect the final result. Unfortunately, today’s reality is different: a change in the label definition, such as adding an article as shown in the figure, can lead to minor prompt variations with drastic changes in the final prediction. As a result, developers remain unaware of the malfunction’s cause and might later abandon the tool due to frustration.
@@
We set out to address these problems by proposing two diagnostic metrics for LLM classifiers that are complementary to accuracy: the sensitivity to the input, which does not depend on the ground truth labels, and the consistency of predictions across examples of the same class. Intuitively, it is desirable to have LLMs that are robust to semantically equivalent variations of the initial prompt, and their predicted labels’ distributions should not vary much across samples of the same class. Striving to improve these two metrics might significantly reduce the unpredictability of LLMs’ behavior in complex software systems running in production; low sensitivity and high consistency might mean that LLMs hallucinate less and have a better understanding of the tasks. Consider the example of Figure 2, taken from our experiments, where the predicted labels’ distribution over samples of a given class completely changes for two semantically equivalent but syntactically different rephrasings of the prompt. This behavior is highly undesirable, and the first step to address it is to define suitable metrics to measure it.
Figure 2: Comparison between predicted class distributions of Llama3 across samples of the same class (TREC dataset, Section 4). A merely syntactic rephrasing of the prompt leads to an abrupt change in the distribution, suggesting low consistency.
We evaluate these metrics on five different datasets, two closed and two open-source LLMs, and three prompting strategies, and we find that they indeed convey different information about the LLMs’ behavior. We use this information to analyze more qualitatively some of the datasets, mimicking the kind of study that a user or an LLM engineer would do when debugging their systems. We argue that these metrics should be included in all automatic prompt engineering/optimization frameworks, and perhaps in the training of LLMs themselves, to progress the development of trustworthy AI systems.
@@ 2 Related Work
It is well-known that variations of the prompt affect stability of the LLM’s accuracy. In a broad study across diverse tasks, McCoy et al. (2023) showed that the performance of the LLM depends on the likelihood of the input prompt and the correct output answer. Such result is consistent with the fact that LLMs are, in essence, autoregressive models trained to maximize a likelihood objective. In light of these considerations, it is unsurprising that simply changing the ordering of the examples in a few-shot prompting strategy can lead to almost random accuracy on sentiment analysis tasks (Zhao et al., 2021). Similar considerations motivated frameworks like FormatSpread, which tries to predict the expected performance under prompt’s variations without accessing the LLM’s weights (Sclar et al., 2024). Also, spurious features in the prompt had severe repercussions on the detection of security vulnerabilities, where LLMs are inconsistent and unfaithful (Ullah et al., 2024). Finally, Yang et al. (2024) study the effect of prompt rephrasings and LLMs’ temperature on classification and uncertainty metrics.
Note that the solution to the spurious feature problem might also reside in a more structured approach: Retrieval Augmented Generation systems (Lewis et al., 2020) or Knowledge Graphs-enhanced LLMs (Luo et al., 2024) reduce hallucinations and dependencies on specific prompts.
@@
Complementary to the discussion in this paper is the estimation of the LLM’s uncertainty. Typically, uncertainty is defined over the different answers of the LLM given the same prompt (Press et al., 2024); being generative models of text, LLMs might produce contradicting predictions due to stochasticity in the output response. Several works have already investigated uncertainty: Kadavath et al. (2022) asked the LLM to provide a numerical score of confidence, whereas Chen et al. (2024) argued that aggregating an LLM’s responses via a majority voting mechanism reveals a non-monotonic relationship between the number of LLM calls and system performance. Motivated by the fact that higher uncertainty should imply lower performances, Huang et al. (2024) define a rank calibration error to quantify deviations from the ideal relationship between the two quantities. In a completely different study with about 400 human participants, Kim et al. (2024) provide evidence that allowing an LLM to express uncertainty in the form of natural language reduces the users’ trust in the system. The authors argue that reducing output uncertainty is crucial for LLMs’ successful adoption. Very recently, Yadkori et al. (2024) proposed an information-theoretic metric to distinuish between epistemic uncertainty, which arises from lack of knowledge about the ground truth, and aleatoric uncertainty, which comes from irreducible randomness. They enable the identification of unreliable model outputs and hallucinations without modifying the training process.
While related to our approach, it does not fall within the scope of this paper to further analyze output uncertainty. Because LLMs can also behave deterministically (or close to it) (Yang et al., 2024), which limits the impact of such uncertainty quantification metrics, this work solely focuses on variations of the input prompts. For a comprehensive study of uncertainty quantification techniques on LLMs, we refer the reader to Huang et al. (2023).
@@
Finally, we discuss techniques to optimize prompts so that the LLMs improve their performance on the task. One cause for bad performances is the bias of the LLM towards samples of a class that appears frequently in the training data; for this reason, a model calibration technique has been proposed to make the predictions more uniform across answers (Zhao et al., 2021). The approach requires access to the LLM’s inner workings, which might not be practical for developers. Automated Prompt Engineer (Zhou et al., 2023) optimizes prompt instructions with the help of an LLM. This approach has also been applied to train the LLM itself, to reduce its susceptibility to adversarial attacks and jailbreaks that overcome the safeguards for ethical use of these systems (Zhou et al., 2024). At the same time, there is still much work to do: LLM-based automatic prompt optimizers struggle to identify the true causes of errors, and we should rather focus on an automated behavior optimization paradigm (Ma et al., 2024).
Prompt optimization is already a step ahead compared to the quantification of sensitivity, and we hope that the definition of suitable metrics will inspire new prompting strategies that do not focus on the sole evaluation of task accuracy.
@@ 3 Methodology
We formally introduce the metrics of interest in a bottom-up fashion, namely sensitivity and consistency. We include uncertainty in the discussion to stress the differences between all quantities.
Let us consider a classification task $\tau$ with $C$ possible classes and a data set of $N$ i.i.d. samples $\mathcal{D}={(\bm{x}_{i},y_{i})}_{i\in\{1,\dots,N\}}$, with $\bm{x}$ being the input text to classify and $y\in\{1,\dots,C\}$ the ground truth label. Moreover, we refer to the subset of samples of class $y$ as $\mathcal{D}_{y}$. We make the typical conditional independence assumption about the data, that is, $p(y,\bm{x})=p(y|\bm{x})p(\bm{x})$. Without loss of generality, we introduce the distribution $p_{\rho_{0}}(\rho)$ defined over all prompts that are semantically equivalent to a reference prompt $\rho_{0}$, which is typically prepared by the user based on task $\tau$; when clear from the context, we will omit the subscript from the notation. Because this distribution depends on the task but is independent of the input distribution, we shall write $p(y,\bm{x},\rho)=p(y|\bm{x},\rho)p(\bm{x})p(\rho)$. It is worth noting that LLMs implement the distribution $p(y|\bm{x},\rho)$, hence we can define the average probability that an LLM predicts the class $y$ under different variants of the same prompt as
$p_{\tau}(y|\bm{x})=\mathbb{E}_{\rho\sim p(\cdot)}[p(y|\bm{x},\rho)].$ (1)
In practice, we can approximate Equation 1 by Monte Carlo integration, using $Q$ prompts $\rho_{1},\dots,\rho_{Q}$ semantically equivalent to $\rho_{0}$:
@@
We then define the sensitivity of an LLM to the prompt to reflect how much the LLM prediction varies under the rephrasings of the original prompt.
@@
Given a categorical distribution $p_{\tau}(y|\bm{x})$ defined as in Equation 1, the sensitivity with respect to the input $\bm{x}$ is its entropy
$S_{\tau}(\bm{x})=-\mathbb{E}_{y\sim p_{\tau}(\cdot|\bm{x})}[\ln p_{\tau}(y|\bm{x})],$ (3)
@@
$C_{y}=\mathbb{E}_{\bm{x},\bm{x}^{\prime}}[C_{y}(\bm{x},\bm{x}^{\prime})]\approx\sum_{\bm{x},\bm{x}^{\prime}\in\mathcal{D}_{y}}\frac{C_{y}(\bm{x},\bm{x}^{\prime})}{|\mathcal{D}_{y}|^{2}}.$ (7)
Intuitively, a consistent LLM produces similar distributions $p_{\tau}(\cdot|\bm{x})$ regardless of the sample $\bm{x}$ of class $y$, indicating that fine-tuning the LLM on the task $\tau$ might improve the situation. An inconsistent LLM, instead, behaves unpredictably among samples of the same class, suggesting that the main cause of low performance is the prompt $\rho$
We argue that, in order to avoid bad surprises in production environments where new LLMs have to be tested and replaced quickly, it might be desirable to select LLMs with low sensitivity and high consistency, which means $S_{\tau}\rightarrow 0$ and $C_{y}\rightarrow 1,\forall y\in\{1,\dots,C\}$. Jointly optimizing the prompt (or the LLM!) against accuracy, sensitivity, and consistency, might result in less hallucinations and a better understanding of the tasks’s context.
Consider the difference of sensitivity and consistency with respect to the output uncertainty, defined by Huang et al. (2024) as the entropy of the LLM predictions for fixed input $\bm{x}$ and prompt $\rho$:
@@
It is clear that the output uncertainty is orthogonal to sensitivity, and it treats the LLM as a stochastic predictor. However, in classification tasks it is reasonable to force a deterministic behavior of the LLM, effectively achieving $U(\bm{x},\rho)=0\ \forall\bm{x},\rho$. In contrast, as of today, there is no solution for achieving zero sensitivity with an LLM.
@@ 4 Experiments
The goal of the experiments is to assess sensitivity and consistency of LLMs on different English-based text classification datasets, analyzing the impact of prompt variations and prompting strategies from a different angle compared to previous works, and showing developers how to use these metrics [^1] https://github.com/nec-research/sensitivity-consistency-LLM.. We ran the LLMs on a 4 Tesla V100 server with 32 GBs of memory and 252 GBs of RAM.
We consider Llama-3-70B-Instruct (Touvron et al., 2023) and Mixtral-8x7B-Instruct-v0.1 (Jiang et al., 2024) as the two open-source LLMs available on our servers [^2] Refer to https://huggingface.co/models., as well as GPT3.5-turbo-0125 and GPT-4o-2024-05-13 (Brown et al., 2020) as closed-source models. The temperature is set to zero and the seed is fixed to 42 to obtain quasi-deterministic behavior. Following Zhao et al. (2021), five multiclass classification datasets are used for the comparison: TREC, a 6-class (and 50 subclasses) question-answering task (Voorhees and Tice, 2000) with 500 test samples, CommittmentBank (CB) as a 3-way classification problem (De Marneffe et al., 2019) with 250 test samples, a binary textual entailment problem (RTE, Dagan et al. (2005)) with 2490 test samples, and the 14-class ontology extraction dataset DBPedia (Zhang et al., 2015) with 2000 balanced test samples. In addition, we also consider the 7-class Web of Science 46985 (WoS) dataset with 2000 balanced test samples (Kowsari et al., 2018). In the event the LLM cannot produce a valid class, we add an extra class label N/A.
We consider three different prompting strategies: a simple strategy, where the prompt consists of the task description and the list of classes; a detail strategy, where we provide a detailed description of each class; and a 1-shot strategy in which, compared to simple, we also provide one example for each class taken from samples that do not belong to the test set. To build different rephrasings of the task description only, we use the aforementioned LLMs. In particular, the prompt asks to rephrase the task description by changing the length or adding unnecessary words as long as the meaning remains the same. As done in Yang et al. (2024), we choose $Q=10$. In the quantitative analyses, we display the average sensitivity and average consistency across samples and classes, as well as the micro F1-score across all $\rho$ for completeness. Since the distributions of values are far from being Gaussian, we do not report standard deviations and refer to more qualitative analyses in later sections that show deviations from the mean values.

### v2

@@ Abstract
Large Language Models (LLMs) changed the way we design and interact with software systems. Their ability to process and extract information from text has drastically improved productivity in a number of routine tasks. Developers that want to include these models in their software stack, however, face a dreadful challenge: debugging LLMs’ inconsistent behavior across minor variations of the prompt. We therefore introduce two metrics for classification tasks, namely sensitivity and consistency, which are complementary to task performance. First, sensitivity measures changes of predictions across rephrasings of the prompt, and does not require access to ground truth labels. Instead, consistency measures how predictions vary across rephrasings for elements of the same class. We perform an empirical comparison of these metrics on text classification tasks, using them as guideline for understanding failure modes of the LLM. Our hope is that sensitivity and consistency will be helpful to guide prompt engineering and obtain LLMs that balance robustness with performance.
@@ 1 Introduction
This famous quote refers to the innate tendency of computer scientists to choose poor names for a program’s variables despite the existence of good coding practices (McConnell, 2004). Practically speaking, this is not a problem as long as the program does its job, but can we still argue the same in the era of Large Language Models (LLMs)?
@@
Figure 1: Example of GPT3.5 behavior when classifying a question in terms of what it is referring to. A slight change in the definition of the class “ENTY” causes a minor prompt variation that disrupts the LLM’s prediction. This happens under the hood, making it very hard for a developer to debug the program. Note that the same might happen, for instance, if the ordering or naming of variables is changed (hence the quote of Section 1).
From a practical standpoint, integrating LLMs in a software infrastructure introduces additional complexities, from choosing good prompt engineering strategies to parsing and controlling the output format of responses. In an effort to simplify the LLM into a straightforward function call, new software libraries like Instructor (Liu, 2024) have emerged, reducing even further the entry barrier of LLMs for the ordinary programmer. Instructor requires to define target labels as static fields of a Python class, which is then automatically converted into a standardized JSON schema. The schema and the input sentence are then used to produce an LLM response via a mechanism known as Function Calling (Kang et al., 2023). Figure 1 provides an example of text classification using Instructor. Here, the assumption of the programmer should be that small changes to the labels’ descriptions, as well as the labels’ ordering in the code, should not affect the final result. Unfortunately, reality is different: a change in the label definition, such as adding an article as shown in the figure, or in variables’ naming can lead to minor prompt variations with drastic changes in the final prediction. As a result, developers remain unaware of the malfunction’s cause and might later abandon the tool due to frustration.
@@
We set out to address these problems by proposing two diagnostic metrics for black-box classifiers that are complementary to accuracy: the sensitivity to the input, which does not depend on the ground truth labels, and the consistency of predictions across examples of the same class. Intuitively, it is desirable to have LLMs that are robust to semantically equivalent variations of the initial prompt, and their predicted labels’ distributions should not vary much across samples of the same class. Striving to improve these two metrics towards low sensitivity and high consistency might significantly reduce the unpredictability of LLMs’ behavior in complex software systems running in production. Consider the illustrative example of Figure 2, where three samples of the same class Person are classified (in probability) very differently when testing different variations of the same prompt. Samples 1 and 3 have lower sensitivity than sample 2, as their distributions are quite stable on the correct class. In addition, sample 2 has low consistency compared to the other samples, which combined with sensitivity indicates “hardness” of classification. The behavior of sample 2 is highly undesirable, and the first step to address it is to formalize metrics that reflect these intuitions.
Figure 2: Predicted class distributions over prompt rephrasings $p_{\tau}$ across three samples of the same class Person (TREC dataset, Section 4). Merely syntactic prompt rephrasings can produce very diverse distributions. For instance, sample 2 is characterized by high sensitivity and (compared to others) low consistency.
We empirically test these metrics, finding that they indeed convey different information about the LLMs’ behavior. We use this information to qualitatively analyze some of the datasets, mimicking the kind of study that an LLM engineer would do when debugging a system. We also show how these metrics help to improve prompts by identifying different sets of problematic samples.
@@ 2 Related Work
It is well-known that variations of the prompt affect stability of the LLM’s accuracy. In a broad study across diverse tasks, McCoy et al. (2023) showed that LLM performances depend on the likelihood of the input prompt and of the correct output answer. Such result is consistent with LLMs’ autoregressive nature, that is, models trained to maximize a likelihood objective. In light of these considerations, it is not surprising that changing the ordering of the examples in a few-shot prompting strategy can lead to almost random accuracy on sentiment analysis tasks (Zhao et al., 2021), or that the nature of prompts affects LLMs’ benchmarks (Ailem et al., 2024). Similar considerations motivated frameworks like FormatSpread, which predicts the expected performance under prompt’s variations without accessing the LLM’s weights (Sclar et al., 2024). Also, spurious features in the prompt have severe repercussions on security vulnerabilities’ detection, where LLMs appear inconsistent and unfaithful (Ullah et al., 2024), while Yang et al. (2024) study the effect of prompt rephrasings and LLMs’ temperature on classification and uncertainty metrics. Finally, we mention the comprehensive benchmark of Liang et al. (2023), where metrics such as invariance to semantic-preserving transformations are computed w.r.t. accuracy changes. Note that the solution to the spurious feature problem might also reside in a more structured approach: Retrieval Augmented Generation (Lewis et al., 2020) or Knowledge Graphs-enhanced LLMs (Luo et al., 2024) reduce hallucinations and prompt dependencies.
@@
Complementary to the discussion in this paper is the estimation of the LLM’s uncertainty. Typically, uncertainty is defined over the different answers of the LLM given the same prompt (Press et al., 2024); being generative models of text, LLMs might produce different predictions due to stochasticity in the output response. Several works have already investigated uncertainty: Kadavath et al. (2022) ask the LLM to provide a score of its confidence, whereas Chen et al. (2024) argued that a majority voting mechanism reveals a non-monotonic relationship between the number of LLM calls and system performance. Motivated by the fact that higher uncertainty should imply lower performances, Huang et al. (2024) define a rank calibration error to quantify deviations from the ideal relationship between the two quantities. In a human study with about 400 participants, Kim et al. (2024) provide evidence that an LLM expressing uncertainty in natural language reduces the users’ trust in the system. Very recently, Yadkori et al. (2024) proposed an information-theoretic metric to distinguish between epistemic and aleatoric uncertainty. They enable the identification of unreliable model outputs and hallucinations without altering the training process.
It does not fall within the scope of this paper to analyze output uncertainty. Because LLMs can also behave pseudo-deterministically (Yang et al., 2024), this limits the impact of some probability-based uncertainty quantification metrics. For a comprehensive study of uncertainty quantification techniques on LLMs, such as verbalization and generation likelihood, we refer the reader to Huang et al. (2023) and Geng et al. (2024).
@@
To improve LLMs performance, one can carefully engineer their prompt. One cause for LLMs’ bad performances is the bias towards over-represented classes; for this reason, Zhao et al. (2021) propose a calibration technique that makes the predictions more uniform across classes. The approach requires access to the LLM’s inner workings, which might be impractical. Automated Prompt Engineer (Zhou et al., 2023) optimizes prompts with the help of an LLM, reducing its susceptibility to adversarial attacks and jailbreaks that overcome the safeguards for ethical use of these systems (Zhou et al., 2024). At the same time, there is still much work to do: LLM-based automatic prompt optimizers struggle to identify the true causes of errors, and we should rather focus on an automated behavior optimization paradigm (Ma et al., 2024). Prompt optimization is already a step ahead compared to the quantification of our metrics, and we hope that this work will inspire new prompting strategies that do not focus on the sole evaluation of task accuracy.
@@ 3 Methodology
We formally introduce the metrics of interest, sensitivity and consistency, in a bottom-up fashion.
Let us consider a classification task $\tau$ with $C$ possible classes and a data set of $N$ i.i.d. samples $\mathcal{D}={(\bm{x}_{i},y_{i})}_{i\in\{1,\dots,N\}}$, with $\bm{x}$ being the input text to classify and $y\in\{1,\dots,C\}$ the ground truth label. Moreover, we refer to the subset of samples of class $y$ as $\mathcal{D}_{y}$. We make the typical conditional independence assumption about the data, that is, $p(y,\bm{x})=p(y|\bm{x})p(\bm{x})$. Without loss of generality, we introduce the distribution $p_{\rho_{0}}(\rho)$ defined over all prompts that are semantically equivalent to a reference prompt $\rho_{0}$, which is typically prepared by the user based on task $\tau$; when clear from the context, we will omit the subscript from the notation. Because this distribution depends on the task but is independent of the input distribution, we shall write $p(y,\bm{x},\rho)=p(y|\bm{x},\rho)p(\bm{x})p(\rho)$. Please note that the distribution $p(\rho)$ is not uniform. In practice, LLMs implement the distribution $p(y|\bm{x},\rho)$, hence we can define the average probability that an LLM predicts the class $y$ under different variants of the same prompt as
$p_{\tau}(y|\bm{x})=\mathbb{E}_{\rho\sim p(\cdot)}[p(y|\bm{x},\rho)],$ (1)
[…]
@@
In this work, the sampler $\mathcal{S}$ is an LLM that is tasked to rephrase prompts (Appendix D shows examples), but it could also be an algorithm that generates text modifications as in Liang et al. (2023). We then define the sensitivity of an LLM to the prompt to reflect how much the LLM prediction varies under the rephrasings of the original prompt.
@@
Given a distribution $p_{\tau}(y|\bm{x})$ defined as in Equation 1, the sensitivity with respect to $\bm{x}$ is the normalized entropy
$S_{\tau}(\bm{x})=-\mathbb{E}_{y\sim p_{\tau}(\cdot|\bm{x})}[\ln p_{\tau}(y|\bm{x})]/\ln(C),$ (3)
@@
$C_{y}=\mathbb{E}[C_{y}(\bm{x},\bm{x}^{\prime})]\approx\sum_{\bm{x},\bm{x}^{\prime}\in\mathcal{D}_{y}}\frac{C_{y}(\bm{x},\bm{x}^{\prime})}{|\mathcal{D}_{y}|^{2}}.$ (7)
Intuitively, a consistent LLM produces similar distributions $p_{\tau}(\cdot|\bm{x})$ regardless of the sample $\bm{x}$ of class $y$. When sensitivity is not 0, being consistent suggests that prompt rephrasings cause similar mistakes across all samples of class $y$, hence a careful tuning of the prompt is required. Instead, an inconsistent LLM behaves unpredictably among samples of the same class, where the same prompt rephrasings cause different mistakes; this might indicate that the problem is not the prompt, rather the classifier itself. Note that, when sensitivity is 0, (in)consistency is uniquely determined by the inherent difficulty of the classification task, as it counts pair-wise class mismatches in the LLM predictions. In this case, consistency will correlate more with classification metrics such as accuracy, thus we expect it to be most useful when sensitivity is high enough, e.g., $0.1$.
We argue that, in order to avoid bad surprises in production environments where new LLMs have to be tested and replaced quickly, it might be desirable to select LLMs with low sensitivity and high consistency, which means $S_{\tau}\rightarrow 0$ and $C_{y}\rightarrow 1,\forall y\in\{1,\dots,C\}$.
Finally, consider how probability-based output uncertainty is defined by Huang et al. (2024) for fixed input $\bm{x}$ and prompt $\rho$:
@@
Output uncertainty is orthogonal to sensitivity, as it does not focus on variations of the prompt $\rho$.
@@ 4 Experiments
The goal of the experiments is to assess how sensitivity and consistency can help in the context of English-based text classification datasets, analyzing the impact of prompt variations and prompting strategies from a different angle compared to previous works, but most importantly showing developers how to use these metrics [^2] The code to reproduce our results is provided in the supplementary material and will be open-sourced at publication.. Therefore, it is outside the scope of this work to perform an extensive benchmark of these metrics across many datasets and LLMs, which would not be helpful to our purpose. large number of LLMs across We ran the LLMs on a 4 Tesla V100 server with 32 GBs of memory and 252 GBs of RAM.
We consider Llama-3-70B-Instruct (Touvron et al., 2023) and Mixtral-8x7B-Instruct-v0.1 (Jiang et al., 2024) as the two open-source LLMs available on our servers, as well as GPT-3.5-turbo-0125 and GPT-4o-2024-08-06 (Brown et al., 2020) as closed-source models. The temperature is set to zero and the seed is fixed to 42 to obtain quasi-deterministic behavior. Following Zhao et al. (2021), five multiclass classification datasets are used for the comparison: TREC, a 6-class (and 50 subclasses) question-answering task (Voorhees and Tice, 2000) with 500 test samples, CommittmentBank (CB) as a 3-way classification problem (De Marneffe et al., 2019) with 250 test samples, a binary textual entailment problem (RTE, Dagan et al. (2005)) with 2490 test samples, and the 14-class ontology extraction dataset DBPedia (Zhang et al., 2015) with 2000 balanced test samples. In addition, we also consider the 7-class Web of Science 46985 (WoS) dataset with 2000 balanced test samples (Kowsari et al., 2018). In the event the LLM cannot produce a valid class, we add an extra class label N/A.
We consider three different prompting strategies: a simple strategy, where the prompt consists of the task description and the list of classes; a detail strategy, where we provide a detailed description of each class; and a 1-shot strategy in which, compared to simple, we also provide one example for each class taken from samples that do not belong to the test set. To build different rephrasings of the task description only, we use the aforementioned LLMs. In particular, the prompt asks to rephrase the task description by changing the length or adding unnecessary words as long as the meaning remains the same. For the purpose of our work, which is only to show how sensitivity and consistency can be used, we fix $Q=30$ after observing that increasing this value did not vary results significantly and was enough for our demonstrative purposes. Also, the “right” value of $Q$ ultimately depends on the user’s compromise between computational resources and stability of resulting statistics. In the quantitative analyses, we display the average sensitivity and average consistency, as well as the micro F1-score across all $\rho$ for completeness. Since the distributions of values are far from being Gaussian, we do not report standard deviations and refer to more qualitative analyses in later sections that show deviations from the mean values.